Fermi National Laboratory


Accelerator Update

Monday May 24
The day shift began with Operations monitoring store 3530, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

Pbar experts began a Debuncher study around 10:45 AM that disrupted stacking.

A Booster transition gate control card faulted at 10:51 AM. Operators managed to reset the card about five minutes later.

At 3:30 PM, Booster technicians turned off a power supply (EAPS) so they could repair a Booster RF station (BRF1). The repair took thirty minutes.

A Pbar expert began a reverse proton study at 4:53 PM.

At 5:22 PM, Booster beam was again stopped to make further repairs on BRF1. Beam resumed about 30 minutes later.

Operations reported at 9:12 PM that a Recycler and a Main Injector (MI) expert had been conducting a combined study for about forty-five minutes. They completed their study at 11:27 PM.

Operations established beam to the Meson MCenter experiment at 11:14 PM.

Operations tried to return Pbar to stacking at 11:35 PM, but a problem with ARF1 delayed this by about fifteen minutes. New ARF1 curves were loaded at this time.

Tuesday May 25
The midnight shift began with Operations monitoring store 3530, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 2:10 AM, a lightning strike caused a Tevatron (TeV) sector C separator to spark.

Operations terminated store 3530 at 5:09 AM.

MCenter finished using beam at 5:11 AM.

Operations put the TeV through a dry squeeze at 5:58 AM.

Operations put the TeV into shot setup at 6:21 AM.

At 7:55 AM, Operations established store 3532 with an initial luminosity of 68.92E30. This set a New Luminosity Record.

Operations began trying to solve a beam problem for the Collider Detector at Fermilab (CDF) at 8:38 AM. After thirty minutes of TeV tuning CDF was able to turn on their detector.

At 9:53 AM, a MI transfer line (A1) had a vacuum pump trip off. Operators kept the A1 beam valves closed for about fifteen minutes while they reset the pumps.

The Recycler began stashing shots of antiprotons at 10:40 AM.

At 10:47 AM, the Booster suffered from a vacuum burst that tripped of several RF stations. They all reset okay.

The Recycler had stashed five transfers of antiprotons, by 11:57 AM. Pbar resumed stacking about ten minutes later.

A MI expert began a slipstacking study at 4:20 PM.

CDF asked Operations at 5:37 PM to fix a TeV beam problem. Operators tuned the beam.

At 7:01 PM, the TeV beam problem came back. Operators adjusted collimators. CDF tuned back on five minutes later.

Wednesday May 26
The midnight shift began with Operations monitoring store 3532 and the Recycler stash, with Pbar stacking, and with MiniBooNE taking beam.

Operations established beam to MCenter at 12:55 AM.

CDF asked Operations to fix a TeV beam problem at 4:08 AM. Operators adjusted the vertical tune.

MCenter finished using beam at 5:52 AM.

Plans for Today
The plans for today include allowing TeV experts to conduct an approximate two-hour end-of-store study and to allow a Tevatron Electron Lens expert to conduct a half-hour study before Operations puts the TeV into the next shot setup. MI experts will use one hour of no beam to conduct a timeline study.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/26/2004   email Fermilab